# From Grunts to Grammar: Emergent Language from Cooperative Foraging

### v1

@@ 6 Discussion
We investigate the language emergence in multi-agent Foraging Games. Through systematic and quantifiable experiments on computational models in simulated environments, we gain valuable insights into key questions in evolutionary linguistics—such as what drives the evolution of language and how it becomes essential for teamwork. Remarkably, without any direct supervision from human languages, agents with limited capacity trained in cooperative settings develop five hallmark properties of human language hockett1960origin . Interchangeability emerges through self-play training or social interactions within larger populations (Table 1). Arbitrariness is demonstrated by two cross-play agents developing non-interchangeable languages that are nonetheless mutually intelligible enough to enable successful cooperation and a high task success rate (Table 1 and 3(a)). Compositionality is indicated by high topsim scores in larger populations, a common metric measuring the structural alignment between messages and environmental semantics (3(a)). Displacement is demonstrated by high decoding accuracy of items’ positions, scores, and spawn times, showing that agents can refer to when and where currently invisible events happen (5(a)). Cultural transmission is evidenced by the gradual degradation of language similarity and success rate across increasing distance in ring-structured populations, despite no direct training between distant agent pairs. (Figure 4, 1(d), and 2(d)). This suggests that culture is transmitted and shaped through interaction within structured social networks. Furthermore, our framework demonstrates that communication can emerge in both implicit and explicit forms, which standard referential games do not support.
Understanding the origins of language is a profound and long-standing challenge that requires insights from linguistics, psychology, neuroscience, and artificial intelligence. Our work represents an initial step toward addressing this complex question. Nonetheless, we recognize several limitations in our current approach. While the training method may mirror aspects of human communication learning, it may not scale effectively to complex environments such as visual games or robotic control tasks. Bridging this gap may require loosening some of the constraints currently imposed, for example, solving more complex tasks may require agents to share gradients or parameters to stabilize the training. Additionally, our framework does not incorporate turn-taking dynamics, which is another challenge that the current methods struggle with. Finally, we evaluate compositionality using a finite set of symbolic inputs. Future work could explore compositional representations that emerge from compressing high-dimensional sensory data into compact latent variables.
Appendix A Learning to Act and Communicate with PPO
@@ Appendix B Theoretical Analysis
Appendix C Additional Figures and Tables
Table S1: Performance of XP agents in the TemporalG environment. $N_{\text{pop}}$: population size. LS: language similarity; Cross-SR: success rate with other agents; Self-SR: success rate in self-play. All values are reported as mean ± standard deviation across held-out agent pairs.
Model $N_{\text{pop}}$ LS Cross-SR Self-SR
XP 3 $0.336\pm 0.012$ $0.975\pm 0.003$ $0.962\pm 0.006$
(a) Fully-Connected: XP
(b) Fully-Connected: XP+SP
(c) Ring-Structured: XP
(d) Ring-Structured: XP+SP
Figure S1: Language Similarity between agent pairs under different social structures and training strategies.
(a) Fully-Connected: XP
(b) Fully-Connected: XP+SP
(c) Ring-Structured: XP
(d) Ring-Structured: XP+SP
Figure S2: Success rate between agent pairs under different social structures and training strategies.
@@ Appendix C Additional Figures and Tables
Figure S3: In t-SNE space, sequences of messages from two different agents are visually separable in the 2 XP Agents setting.
(a) Episodic return
(b) Action entropy
(c) Message entropy
(a) Episodic return
(b) Action entropy
(c) Message entropy
Appendix D Metrics
Topographic Similarity (topsim)
Topographic similarity measures the alignment between the structure of the message space and the semantic space (e.g., environmental states or properties). It is defined as the Spearman rank correlation between all pairwise distances in these two spaces [4, 37].
Language Similarity (LS)
quantifies the token-level similarity between two agents based on their communication over multiple episodes, starting from the same initial condition. For a given episode $e$, the agents $i$ and $j$ produce their respective message sequences $M^{(e)}_{i}$ and $M^{(e)}_{j}$. Let $\mathcal{D}_{\text{edit}}(\cdot)$ represent the normalized edit distance function, and let $N_{e}$ denote the total number of evaluation episodes. The Language Similarity (LS) between agents $i$ and $j$ is defined as:
[…]
Interchangeability (IC)
Appendix E Generalization: Varying Grid Size and Adding Obstacles
Figure S6: Visualization of central obstacles in the environment. Obstacles are shown as gray squares randomly placed in the middle of the map. Agents cannot move through gray squares, which increases coordination difficulty.
We analyze the generalization and robustness of 3 XP agents in both the ScoreG and TemporalG environments by evaluating performance across varying grid sizes and obstacle configurations. While not directly focused on communication, this analysis serves to evaluate how well learned behaviors and coordination strategies transfer to novel spatial configurations.
Generalization to unseen grid sizes.
Tables S2 and S3 report agent performance when evaluated on larger grid sizes than those seen during training. XP agents trained on $5\times 5$ environments experience a sharp performance drop as grid size increases in both tasks. In the ScoreG task, the success rate declines from $0.964$ on the training grid to $0.333$ on a $9\times 9$ grid, with longer episode lengths indicating increased exploration difficulty. A similar trend is observed in TemporalG, where success drops drastically from $0.971$ to $0.179$ between $5\times 5$ and $7\times 7$ grids. These results suggest that generalization to larger spatial layouts remains a challenge.
Effect of central obstacles.
To evaluate the agents’ robustness under spatial constraints, we introduce impassable obstacles in the center of the map (visualized in Figure S6). In the ScoreG game (Table S4), performance degrades progressively as the number of obstacles increases. Success rate falls from $0.960$ with no obstacles to just $0.053$ with four obstacles, while episode length grows, indicating less efficient navigation and coordination. The effect is even more pronounced in TemporalG (Table S5), where communication relies heavily on agents meeting at precise times and locations. With just one obstacle, the success rate drops from $0.971$ to $0.275$, and with two obstacles, it drops to $0.061$.
Table S2: Grid Size Generalization performance in the ScoreG environment. Agents were trained on a $5\times 5$ grid and evaluated on larger, unseen grid sizes without fine-tuning. SR: average success rate; Episode Length: average number of steps per episode.
Grid Size Success Rate (SR) Episode Length
$5\times 5$ $0.964\pm 0.022$ $5.03\pm 0.22$
$6\times 6$ $0.925\pm 0.040$ $9.09\pm 1.40$
$7\times 7$ $0.643\pm 0.191$ $14.34\pm 3.47$
$8\times 8$ $0.446\pm 0.228$ $16.34\pm 4.76$
$9\times 9$ $0.333\pm 0.214$ $17.49\pm 5.44$
Table S3: Grid Size Generalization in the TemporalG environment. XP agents are trained on $5\times 5$ grids and evaluated on larger unseen grid sizes without fine-tuning. SR: average success rate; Episode Length: average number of steps per episode.
Grid Size Success Rate (SR) Episode Length
$5\times 5$ $0.971\pm 0.010$ $15.37\pm 0.10$
$6\times 6$ $0.501\pm 0.080$ $19.48\pm 0.64$
$7\times 7$ $0.179\pm 0.053$ $23.93\pm 0.74$
Table S4: Performance of 3 XP agents in the ScoreG environment with increasing numbers of central obstacles. As obstacle count increases, success rate decreases and episode length increases, indicating reduced coordination efficiency under spatial constraints.
Number of Obstacles Success Rate (SR) Episode Length
0 $0.960\pm 0.026$ $4.98\pm 0.19$
1 $0.528\pm 0.033$ $5.28\pm 0.24$
2 $0.280\pm 0.049$ $5.66\pm 0.36$
3 $0.131\pm 0.047$ $6.16\pm 0.42$
4 $0.053\pm 0.027$ $6.46\pm 0.56$
Table S5: Performance of XP agents in the TemporalG environment with varying numbers of central obstacles. Increased obstacles reduce success rates significantly, suggesting that communication about temporally grounded information becomes less effective when spatial coordination is constrained.
Number of Obstacles Success Rate (SR) Episode Length
0 $0.971\pm 0.010$ $15.37\pm 0.10$
1 $0.275\pm 0.019$ $14.75\pm 0.13$
2 $0.061\pm 0.008$ $14.61\pm 0.37$
Appendix F Ablation: Vocabulary Size
@@
In ScoreG, the agents typically complete the game in approximately 5-6 time steps because we introduce time pressure to encourage agents to finish the game as soon as possible. Varying the vocabulary size alters the capacity of information that can be conveyed through the message sequence. Here we train XP agents with $N_{\text{pop}}=3$ with different vocabulary sizes. The default vocabulary size we used in the main text was 4 (i.e., $\mathbf{m_{t}}\in\{0,1,2,3\}$). A smaller vocabulary should, in principle, encourage greater compositionality due to a smaller capacity. As shown in Table S6, compositionality remains relatively stable across vocabulary sizes of $4$, $8$, and $16$. However, when the vocabulary size increases to $32$, the topsim score drops to $0.25$, indicating a notable decline in compositional structure.
@@
Appendix G Training Details
We train our models a single GeForce RTX 4080. However, due to the small model size and low environment complexity, training can also be run entirely on a CPU as well. For populations smaller than 3, training takes approximately 16 hours, depending on the CPU and the number of available cores. The longest training time occurs with a population size of 15 and takes up to 3 days. All experiments are repeated with three different random seeds. We use the ADAM optimizer [31] with an initial learning rate of $0.00025$, which linearly decays over time. We report the hyperparameters of the architecture and training algorithm below.
Table S7: Neural architecture hyperparameters.
Component Hyperparameter
Visual encoder 4-layer MLP: [256, 256, 128, 16]
Visual input shape $(1,5,5)$ grayscale grid
Position encoder Linear(2, 4)
Message embedding size 16
Message encoder Embedding(Vocab. Size, 16) + Linear(16, 16)
LSTM hidden size 128
LSTM input size $16$ (visual) $+$ $4$ (position) $+$ $16$ (message) = 36
Action head Linear(128, num_actions)
Value head Linear(128, 1)
Message head Linear(128, 16)
Table S8: PPO hyperparameters.
Hyperparameter Value
Total time steps $2\times 10^{9}$
Learning rate $2.5\times 10^{-4}$
Number of environments 128
Number of steps per rollout 32
Number of minibatches 4
Number of update epochs 4
Discount factor $\gamma$ 0.99
GAE $\lambda$ 0.95
Clip coefficient 0.1
Clip value loss True
Normalize advantages True
Action entropy coefficient 0.01
Message entropy coefficient 0.002
Value function coefficient 0.5
Max gradient norm 0.5
Learning rate annealing True
Target KL divergence None
Batch size 4096
Minibatch size 1024
Optimizer Adam ($\epsilon=10^{-5}$)

### v2

@@ 6 Discussion
Understanding the origins of language is a profound and long-standing challenge that requires insights from linguistics, psychology, neuroscience, and artificial intelligence. Our work represents an initial step toward addressing this complex question. We investigate the language emergence in multi-agent Foraging Games. Through systematic experiments on computational models in simulated environments, we gain valuable insights into key questions in evolutionary linguistics—such as what drives the evolution of language and how it becomes essential for teamwork. To further probe this point, we conducted a non-cooperative control experiment (Sec. D.3), which showed that when agents pursue independent goals, communication provides no advantage and does not emerge meaningfully. This reinforces cooperation as a key ecological pressure in the evolution of language. Remarkably, without any direct supervision from human languages, agents with limited capacity trained in cooperative settings develop five hallmark properties of human language hockett1960origin . Interchangeability emerges through self-play training or social interactions within larger populations (Table 1). Arbitrariness is demonstrated by two cross-play agents developing non-interchangeable languages that are nonetheless mutually intelligible enough to enable successful cooperation and a high task success rate (Table 1 and 4(a)). Compositionality is indicated by high topsim scores in larger populations, a common metric measuring the structural alignment between messages and environmental semantics (3(a)). Displacement is demonstrated by high decoding accuracy of items’ positions, scores, and spawn times, showing that agents can refer to when and where currently invisible events happen (5(a)). Cultural transmission is evidenced by the gradual decline in LS and SR with increasing distance in structured populations, despite no direct training between distant pairs (Figure 4, 5(d), 6(d)). This suggests that culture is transmitted and shaped through interaction within structured social networks. Our framework also shows that communication can emerge in both implicit and explicit forms, which standard referential games do not support.
Despite these contributions, our work has several limitations. The language emerging in FG is not directly comparable to natural language, remaining simple and lacking syntax and morphology. FG also offers limited semantic richness, as agents primarily communicate about a narrow set of task variables (time, location, score). Progress toward more human-like emergent language will require semantically richer, open-ended environments that encourage broader, compositional, and syntactically structured communication. Finally, our framework does not model turn-taking dynamics, a further challenge for current approaches.
Appendix A Implementation and Training Details
A.1 Learning to Communicate with PPO
@@
A.2 Model Training
We train our models on a single GeForce RTX 4090. However, due to the small model size and low environment complexity, training can also be run entirely on a CPU as well. For populations smaller than 3, training takes approximately 16 hours, depending on the CPU and the number of available cores. The longest training time occurs with a population size of 15 and takes up to 3 days. All experiments are repeated with three different random seeds. We use the ADAM optimizer [35] with an initial learning rate of $0.00025$, which linearly decays over time. We report the hyperparameters of the architecture and training algorithm in Table S1 and Table S2. The examples of learning dynamics for ScoreG and TemporalG are shown in Figure S1 and Figure S2.
Table S1: Neural architecture hyperparameters.
Component Hyperparameter
Visual encoder 4-layer MLP: [256, 256, 128, 16]
Visual input shape $(1,5,5)$ grayscale grid
Position encoder Linear(2, 4)
Message embedding size 16
Message encoder Embedding(Vocab. Size, 16) + Linear(16, 16)
LSTM hidden size 128
LSTM input size $16$ (visual) ++ 44 (position) ++ $16$ (message) = 36
Action head Linear(128, num_actions)
Value head Linear(128, 1)
Message head Linear(128, 16)
Table S2: PPO hyperparameters.
Hyperparameter Value
Total time steps $2\times 10^{9}$
Learning rate $2.5\times 10^{-4}$
Number of environments 128
Number of steps per rollout 32
Number of minibatches 4
Number of update epochs 4
Discount factor $\gamma$ 0.99
GAE $\lambda$ 0.95
Clip coefficient 0.1
Clip value loss True
Normalize advantages True
Action entropy coefficient 0.01
Message entropy coefficient 0.002
Value function coefficient 0.5
Max gradient norm 0.5
Learning rate annealing True
Target KL divergence None
Batch size 4096
Minibatch size 1024
Optimizer Adam ($\epsilon=10^{-5}$)
(a) Episodic return
(b) Action entropy
(c) Message entropy
(a) Episodic return
(b) Action entropy
(c) Message entropy
A.3 Metrics
Topographic Similarity (topsim)
Topographic similarity measures the alignment between the structure of the message space and the semantic space (e.g., environmental states or properties). It is defined as the Spearman rank correlation between all pairwise distances in these two spaces [4, 41]. A higher topsim indicates a more compositional and consistent mapping between semantic meanings and messages. We define the semantic space using ground-truth attributes of the environment, specifically the scores and positions of items in the ScoreG. To implement the topographic similarity metric, we use the EGG toolkit [33].
Language Similarity (LS)
quantifies the token-level similarity between two agents based on their communication over multiple episodes, starting from the same initial condition. A higher score means the agents tend to use the same messages in similar contexts, indicating stronger convergence in their communication strategies. For a given episode $e$, the agents $i$ and $j$ produce their respective message sequences $M^{(e)}_{i}$ and $M^{(e)}_{j}$. Let $\mathcal{D}_{\text{edit}}(\cdot)$ represent the normalized edit distance function, and let $N_{e}$ denote the total number of evaluation episodes. The Language Similarity (LS) between agents $i$ and $j$ is defined as:
$\text{LS(i,j)}=\frac{1}{N_{e}}\sum_{e=1}^{N_{e}}1-\mathcal{D}_{\text{edit}}(M^{(e)}_{i},M^{(e)}_{j})$ (5)
Interchangeability (IC)
@@ Appendix B Theoretical Analysis
Appendix C Additional Experiments and Results on Cultural Transmission
(a) FC
(b) Ring
(c) Clq
(d) WS
(e) LRC
Figure S3: Social network structures: Nodes represent agents, and edges denote pairs of agents that are co-trained. Network structures are abbreviated as follows: FC – Fully Connected, Ring – Ring Structure, Clq – Ring with Cliques, WS – Watts-Strogatz, and LRC – Ring with Long-Range Connections
@@ Appendix C Additional Experiments and Results on Cultural Transmission
Figure S4: Message embeddings under different training regimes. In the t-SNE space of embedding vectors, sequences of messages from two agents are clearly separable in the 2-XP-Agents setting. (a) Messages from different agents occupy distinct regions of the embedding space (indicating different languages). (b-c) Messages from different agents largely overlap (indicating similar languages).
(a) FC: XP
(b) FC: XP+SP
(c) Ring: XP
(d) Ring: XP+SP
Figure S5: Language Similarity (LS) between agent pairs under different social structures and training strategies. (a–b) Under the fully connected (FC) social network, both training regimes (XP and XP+SP) yield similar languages across agents. (c) With XP training in the Ring network, agents develop different languages from their partners (evident from the color switching between light and dark along the axis perpendicular to the diagonal). (d) With XP+SP training in the Ring network, agents and their partners converge to similar languages.
(a) FC: XP
(b) FC: XP+SP
(c) Ring: XP
(d) Ring: XP+SP
Figure S6: Success Rates (SR) between agent pairs under different social structures and training strategies. (a–b) Both XP and XP+SP produce homogeneous SR across all pairs. (c) With XP training in the Ring network, agents fail when paired with copies of themselves. (d) With XP+SP training in the Ring network, agents succeed when paired with themselves.
(a) LS
(b) SR
Figure S7: More cliques and long-range connections both enhance cultural transmission ($N_{\text{pop}}=15$): We plot LS and SR (See Sec. A.3) as a function of the shortest-path distance between two agents. A distance of 1 indicates that two agents were co-trained, while a distance greater than 1 indicates that the agents were never paired during training. Note that the maximum shortest path of LRC is only 4 while other cases are 5.
We further examine whether adding additional connections, while keeping social networks sparse, can enhance cultural transmission. To this end, we employ two networks with small-world properties, as illustrated in Figure S3. The first is the Watts-Strogatz (WS) network, constructed with parameters $k=4$ (neighbors) and $\beta=0.2$. The second is a ring structure with long-range connections (LRC) network, built upon a standard ring lattice with additional long-range edges. Specifically, the LRC network is generated by iteratively linking the farthest nodes on the ring lattice, serving as a simple proxy for reducing average path length and thereby producing a more small-world-like structure.
More cliques and long-range connections both enhance cultural transmission [ScoreG-P15-XP] We examine the impact of adding cliques or long-range connections to the ring-structured network on language transmission. Results are shown in 7(a) and 7(b). For WS and LRC, both LS and SR decrease slightly with social distance, demonstrating successful language transmission to non-co-trained agents in the population. For Clq, LS at distance 1 is already lower and declines more steeply with distance than the others. SR also falls sharply, dropping from $0.95$ at distance $1$ to $0.55$ at distance $5$. This result is intuitive because Clq is constructed by adding only five connections to the ring network, which may not be enough for the population to form similar and compatible languages, whereas the others are constructed by adding many more connections to the ring network.
Appendix D ScoreG: Additional Results and Analyses
D.1 Ablation: Disembodiment (Referential Game)
Figure S8: Decoding item states from messages produced in a no-body referential game. V-Pos/H-Pos are item positions, and Score is item value. Integer Msg uses raw message chains composed of integer sequences. Msg Embedding uses chained embeddings mapped from a lookup table. Error bars show standard deviations.
In this variant, agents are immobilized and remain fixed near their respective items. The game structure and goal are the same as those in ScoreG, but without requiring agents to move to pick up the target item. Specifically, the action space is restricted to {select my item, select another item, idle}. The game succeeds only if both agents correctly select the target item at the final timestep ($t=6$). Any premature selection causes immediate failure. Rewards are binary: +1+1 for success and 1-1 for failure, with no advantage for finishing early since the episode length is fixed.
As shown in Figure S8, agents exchange messages conveying item score information, but not spatial information, a result confirmed by chance-level decoding accuracy. Each agent still observes its own static position, which in principle allows item locations to be inferred. However, this positional information is task-irrelevant, as agents cannot move.
D.2 Ablation: Unidirectional Communication
We study a setting where only one agent (the sender) can transmit messages, while the other (the receiver) cannot send messages back—i.e., the sender receives no information from its partner. Two of the three agents are randomly selected and assigned as sender or receiver, and XP training is used. This setup reflects the unidirectionality typical of referential games.
Agents are evaluated under two conditions: Normal (scores sampled from $2,4,6,\dots,248]$) and Hard (scores sampled from $160,162,\dots,240]$). The result is shown in Table S3. In the Normal condition, agents achieve a high SR of $0.829$ because certain score configurations allow effective guessing. For instance, if the sender observes a low score of 32 and the receiver observes 232, the receiver infers that the sender sees 32 and can wait, while the sender correctly judges that its item is unlikely to be a target and searches for the other.
In the Hard condition, however, SR drops to chance level ($0.508$). Because all scores are moderate, agents cannot reliably infer whether their item is a target. This result underscores the importance of bidirectional communication for achieving high SR in the ScoreG game.
Table S3: Performance of three XP agents with unidirectional communication under different test conditions.
Test Condition SR Length
Normal $0.829\pm 0.022$ $5.166\pm 0.162$
Hard $0.508\pm 0.011$ $5.934\pm 0.503$
D.3 Ablation: No Cooperation
Table S4: Ablation study on non-cooperative tasks (IndividualG). We report average rewards for two agents and the average episode length (mean $\pm$ standard deviation) under three communication conditions.
Condition Avg. Reward (First Player) Avg. Reward (Second Player) Avg. Length
Normal $1.770\pm 0.180$ $1.878\pm 0.194$ $13.762\pm 1.240$
Ablate-Noise $1.822\pm 0.152$ $1.910\pm 0.171$ $13.300\pm 1.364$
Ablate-Zero $1.740\pm 0.135$ $1.945\pm 0.179$ $13.148\pm 1.124$
A leading hypothesis suggests that language, or communication more broadly, emerges because agents need to share their intentions or knowledge to cooperate effectively. To test this hypothesis, we design an additional experiment in which agents are embodied in the same environment but pursue individual goals (i.e., each agent separately picks up any items in the map). In this setting, there is no mutual incentive for agents to pick up the same item—in other words, shared intentionality is removed. We use the same neural network architecture and training regime (XP) as in the ScoreG experiments. Agents can still send and receive messages from their partners, but we hypothesize that communication will be less useful because coordination is not required to achieve shared rewards. We describe this environment, called IndividualG, below.
IndividualG
The game structure is similar to ScoreG, with agents placed in a $6\times 6$ grid world and sharing the same action space. However, there are several key differences. First, no scores are assigned to items, and each item requires only a single agent to pick it up. Consequently, no shared reward is available to cooperative agents, and communication is unnecessary to succeed. Second, an agent receives a reward of +1 for each item it successfully collects. Items cannot be picked up by two agents simultaneously. There are four items in total, so the maximum reward for a single agent is 4. The Nash equilibrium is for each agent to pick up two items. The game ends after 20 time steps or once all items have been collected.
Results
The goal of this experiment is to test whether communication is useful in non-cooperative tasks. In other words, can meaningful communication still emerge when cooperation is unnecessary? We train three XP agents and evaluate them under three conditions: Normal, Ablate-Zero, and Ablate-Noise. In Ablate-Zero, agents’ messages are replaced with zero tokens, preventing the transmission of meaningful signals. In Ablate-Noise, agents’ messages are replaced with random samples drawn from a uniform distribution instead of the policy’s output distribution. This setup forces agents to misinterpret their partners’ messages if the communication learned during training had any effect. In each episode, two of the three agents are randomly assigned as the First Player and Second Player. We report the average reward (corresponding to the average number of items collected per episode) for both players to test for potential asymmetries in their behavior. Results are shown in Table S4. Across all three conditions, agents achieve similar performance, with no observable difference between the first and second players. This suggests that communication does not provide a measurable advantage when agents pursue independent rewards. In other words, without shared goals or incentives for coordination, the environment does not pressure agents to develop or rely on meaningful communication.
D.4 Analysis: Varying Vocabulary Size
@@
In ScoreG, the agents typically complete the game in approximately 5-6 time steps because we introduce time pressure to encourage agents to finish the game as soon as possible. Varying the vocabulary size alters the capacity of information that can be conveyed through the message sequence. Here we train XP agents with $N_{\text{pop}}=3$ with different vocabulary sizes. The default vocabulary size we used in the main text was 4 (i.e., $\mathbf{m_{t}}\in\{0,1,2,3\}$). A smaller vocabulary should, in principle, encourage greater compositionality due to a smaller capacity. As shown in Table S5, compositionality remains relatively stable across vocabulary sizes of $4$, 88, and $16$. However, when the vocabulary size increases to $32$, the topsim score drops to $0.25$, indicating a notable decline in compositional structure.
@@
D.5 Generalization to Unseen Positions
We study how well communication generalizes when it comes to unseen object positions. We trained three XP agents to perform the ScoreG task on 17 food locations in a 5×5 grid and evaluated their generalization to 8 unseen locations. Train and test food locations are randomly selected. The results are shown in Table S6. While agents with communication perform well with SR of $0.953$ on seen locations, SR drops to $0.756$ when agents are tasked to pick up objects at the unseen locations. We further ask whether the observed generalization gap arises from communication itself or from limitations of the underlying RL method. To test this, we train three XP agents that can observe all scores and each other but cannot communicate—essentially a standard MARL setup. The agents are trained on 17 item locations and evaluated on 8 unseen locations. As shown in Table S7, SR drops from $0.981$ to $0.823$, indicating that the gap stems from the inherent limits of the MARL algorithm rather than communication alone. Developing novel methods to improve the generalization of multi-agent coordination under unseen spatial configurations is an important direction for both MARL in general and emergent communication in embodied settings. However, this lies beyond the scope of the present study and we leave it for future work.
Table S6: Performance of three XP agents with communication on seen vs unseen locations.
Location SR Length
Seen $0.953\pm 0.017$ $5.197\pm 0.099$
Unseen $0.756\pm 0.022$ $5.154\pm 0.229$
Table S7: Performance of three XP agents without communication on seen vs unseen locations.
Location SR Length
Seen $0.981\pm 0.006$ $4.389\pm 0.038$
Unseen $0.823\pm 0.023$ $4.092\pm 0.062$
D.6 Analysis of Below-Chance Success Rates in Implicit Communication
In Inv-NoCom setting, the success rate (SR) falls below the chance level (See 5(c)), which indicates a systematic coordination failure rather than random performance. To better understand this phenomenon, we conduct two analyses: (i) ablation of the speed reward, and (ii) categorization of failure outcomes.
Ablation of the speed reward.
The speed reward encourages efficient task completion. Removing this reward does not substantially change SR, which remains below chance. However, the absence of the speed reward increases the average episode length (Table S8). This result shows that the speed reward primarily affects efficiency but does not resolve the core coordination issue that leads to low SR.
Table S8: Effect of the speed reward on success rate (SR) and average episode length.
Reward Condition SR Length
With speed reward 0.438 5.607
Without speed reward 0.446 7.950
Failure outcome statistics.
To diagnose why agents fail, we categorize episode outcomes into three groups: correct pickups, wrong pickups, and no pickups. Table S9 shows the distribution. Failures are dominated by wrong pickups (34.8%), followed by cases where agents never pick up the same item (27.4%). Together, these account for the below-chance success rate and show that agents struggle to coordinate reliably without communication.
Table S9: Distribution of episode outcomes without the speed reward.
Outcome Percentage
Success: Correct Pickup 37.8%
Failed: Wrong Pickup 34.8%
Failed: No Pickup 27.4%
Overall, these findings show that below-chance SR in Inv-NoCom setting does not result from the speed reward but from miscoordination, most prominently through wrong item pickups.
D.7 ScoreG: Bumping and Cell Occupancy
Each grid cell can be occupied by at most one entity (either an agent or an item) at any given time. If an agent attempts to move into a cell that is already occupied by another agent, the move is blocked and the agent remains in its original position. We refer to such blocked moves as bumping events.
Although bumping could theoretically provide a minimal form of implicit communication (e.g., signaling another agent’s location), it constitutes only a very low-bandwidth channel and becomes difficult to coordinate once agents are spatially separated. In practice, bumping is rare and therefore does not play a substantive role in coordination or communication. Table S10 summarizes the average number of bumps observed per episode across different conditions.
Table S10: Average bumps per episode.
Condition Success Failed
Inv-Com 0.179 0.065
Inv-NoCom 0.043 0.057
Vis-NoCom 0.015 0.020
These results confirm that while bumping exists as an artifact of the grid-world dynamics, it is too infrequent (<1<1 bump per episode) to support reliable strategies or emergent communication.
Appendix E TemporalG: Additional Results and Analyses
Table S11: Performance of XP agents in the $5\times 5$ TemporalG environment with the fixed duration of 6. $N_{\text{pop}}$: population size. LS: language similarity; Cross-SR: success rate with other agents; Self-SR: success rate in self-play. All values are reported as mean ± standard deviation across held-out agent pairs.
Model $N_{\text{pop}}$ LS Cross-SR Self-SR
XP 3 $0.336\pm 0.012$ $0.975\pm 0.003$ $0.962\pm 0.006$
E.1 Implicit Communication
As shown in Table S12, three XP agents trained without communication under the invisible-partner condition (Inv-NoCom) still achieved relatively high success at a fixed duration of 6, likely by adopting a waiting strategy—that is, delaying movement to infer the spawn order. Their longer episode lengths, compared to the other two conditions (Inv-Com and Vis-NoCom), support this interpretation. However, when the fixed duration was increased, the number of possible spawn times doubled (from 6 to 12), rendering the waiting strategy ineffective and causing the success rate of Inv-NoCom agents to fall below chance level. Similar to what was observed in ScoreG experiment in section 5, Vis-NoCom agents achieved performance comparable to Inv-Com, suggesting that they exploited visual cues for implicit coordination, especially under the limited set of six possible spawn times. In the Inv-Com-Noise condition, agents were trained with explicit communication but tested without it (messages were always zero). The low SR observed in this condition implies that meaningful messages are necessary for agents trained with communication (Inv-Com) to succeed at the task.
Table S12: Implicit communication in the $8\times 8$ TemporalG with fixed durations of 6 and 12. Length denotes the average length of successful episodes (not applicable in the bottom-right cell, where some runs yielded zero successes). Abbreviations are the same as those reported in Figure 6.
Condition Fixed Duration = 6 Fixed Duration = 12
SR Length SR Length
Inv-Com $0.980\pm 0.035$ $23.990\pm 0.618$ $0.961\pm 0.021$ $30.023\pm 0.775$
Inv-NoCom $0.728\pm 0.155$ $28.716\pm 2.341$ $0.419\pm 0.137$ $31.738\pm 1.780$
Vis-NoCom $0.939\pm 0.093$ $24.516\pm 0.307$ $0.862\pm 0.110$ $30.546\pm 0.406$
Inv-Com-Noise $0.313\pm 0.040$ $28.666\pm 0.774$ $0.003\pm 0.001$
E.2 Rendezvous Behavior Analysis
Figure S9: Rendezvous maps of first meeting points across conditions. Values represent normalized frequencies of encounters on the $5\times 5$ grid. Inv-Com and Vis-NoCom agents consistently meet and converge near the center, while Inv-NoCom agents rarely meet and show no strong central bias.
Restricting the communication range forces agents to meet in physical proximity before they can exchange information. This environmental constraint gives rise to rendezvous points: locations where agents repeatedly converge to initiate communication. Despite this restriction, displacement remains necessary. Agents must explore the grid to observe items, store these observations in memory, and then communicate the information once they encounter their partner. Since neither agent can directly infer when all items have appeared without exchanging messages, coordination still relies on communication.
To analyze where meetings occur, we compute a rendezvous map that records the normalized frequency of the first meeting location across episodes. The resulting heatmaps reveal a pronounced central bias: agents overwhelmingly converge near the middle of the $5\times 5$ grid (See Figure S9). This bias emerges because the center minimizes travel distance and maximizes the probability of encounter under partial observability.
When comparing conditions, clear differences emerge. Inv-Com agents always meet (100% of episodes) and almost exclusively converge in the central cells, reflecting highly coordinated rendezvous behavior. Vis-NoCom agents also meet in every episode, showing a similar central bias, suggesting that visibility supports coordination even without explicit messages. By contrast, Inv-NoCom agents succeed in meeting in only 9.9% of episodes, and when they do meet, the rendezvous points are weakly defined and scattered across the grid.